OpenStudy (anonymous):

Fractional indices, please answer the following questions while showing all the steps. thank you. Find the value of x when: 1) 4^x = 64 2) x^2 = 4/9 3) 8^x = 2

OpenStudy (anonymous):

\[4^x=4^3\]

OpenStudy (anonymous):

therefore wth same base x ll be 3

OpenStudy (anonymous):

\[2)x^2=4/9\]

OpenStudy (anonymous):

ll be multiply by 9 both side

OpenStudy (anonymous):

\[9(x^2)=(4/9)9\]

OpenStudy (anonymous):

\[9x^2=4\]

OpenStudy (anonymous):

\[(3x)^2=2^2\]

OpenStudy (anonymous):

therefore same power ll be

OpenStudy (anonymous):

\[3x=2\]

OpenStudy (anonymous):

Could you please answer #1 again, i did not understand

OpenStudy (anonymous):

divide by 3 both side

OpenStudy (anonymous):

\[x=2/3\]

OpenStudy (anonymous):

but how do you know to divide by 3?

OpenStudy (anonymous):

coz we find the value of x if it like this 3x=2 u should divide by 3

OpenStudy (anonymous):

let me show you an example of another question. and using that help me answer these questions okay.

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

here is an example: 16^1/4 = 4sqrrt16 = 2

OpenStudy (anonymous):

thats the method we are using

OpenStudy (anonymous):

one sec let me show you a drawing

OpenStudy (anonymous):

u mean \[16^1/4=4\sqrt{16}=2\]

OpenStudy (anonymous):